### v2

# Segment Anything in 3D Gaussians

Abstract
3D Gaussian Splatting has emerged as an alternative 3D representation of Neural Radiance Fields (NeRFs), benefiting from its high-quality rendering results and real-time rendering speed. Considering the 3D Gaussian representation remains unparsed, it is necessary first to execute object segmentation within this domain. Subsequently, scene editing and collision detection can be performed, proving vital to a multitude of applications, such as virtual reality (VR), augmented reality (AR), game/movie production, etc. In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters. We refer to the proposed method as SA-GS, for Segment Anything in 3D Gaussians. Given a set of clicked points in a single input view, SA-GS can generalize SAM to achieve 3D consistent segmentation via the proposed multi-view mask generation and view-wise label assignment methods. We also propose a cross-view label-voting approach to assign labels from different views. In addition, in order to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme. Extensive experiments demonstrate that SA-GS achieves high-quality 3D segmentation results, which can also be easily applied for scene editing and collision detection tasks. Codes will be released soon.
1 Introduction
Figure 1: We propose an effective 3D object segmentation method in 3D Gaussian without any training process. The segmented mask contributes to the realization of various applications, including collision detection, translation, rotation and removal, etc.
3D scene understanding is a challenging and crucial task in computer vision and computer graphics, which involves scene reconstruction from images or videos and the perception of a given 3D real-world environment. Researchers have conducted extensive studies in scene reconstruction and 3D scene perception in recent years. For instance, Neural Radiance Fields (NeRF) [20, 4, 10, 28] have significantly contributed to the progress of 3D scene reconstruction by representing scenes in an implicit way. In the field of scene perception, continuous research is being conducted on 3D detection and semantic segmentation, based on the representation of range images [18, 6], point clouds [14, 7, 24, 27, 9], and Bird’s Eye View (BEV) [15, 31, 23, 22, 25]. Although current methods have attained noteworthy success in 3D scene understanding, the time-consuming nature of NeRF and the high costs associated with 3D data collection pose challenges for scaling up these approaches.
Alternatively, SA3D  [3] addresses the aforementioned challenges by enhancing the 2D foundational model (i.e. SAM  [13]) with 3D perception through the incorporation of a 3D representation model. SA3D demonstrates notable success by leveraging NeRF to generate 2D views and aligning 2D segmentation results with the 3D scene. However, the substantial training and rendering time costs associated with NeRF make it impractical for large-scale 3D scenes. Consequently, in this work, we focus on 3D Gaussian Splatting (3D-GS) [11] as a prospective method for modeling static 3D scenes. 3D-GS characterizes intricate scenes by employing numerous colored 3D Gaussians, rendering them into camera views through splatting-based rasterization. Through differentiable rendering and gradient-based optimization, the positions, sizes, rotations, colors, and opacities of these Gaussians can be finely tuned to accurately represent the 3D scene, availing for the comprehension of a 3D environment.
Benefiting from the properties of high-quality and real-time rendering in 3D-GS, it is impressive to use these explicit representations to capture the shape and appearance of complex objects. Specifically, based on the 3D Gaussian volume, we can effectively achieve scene perception via equipping the 2D foundation model as SA3D  [3] statement. Different from SA3D leveraging implicit NeRF as a 3D prior to connecting multiple 2D views, our method (SA-GS) applies effect and explicit 3D-GS representation to perform 3D object segmentation. Subsequently, scene editing and collision detection can be executed readily, as shown in Figure 1, which is essential to various applications, such as virtual reality(VR), augmented reality(AR), and games/movie production. Moreover, in contrast to the existing labor-intensive data collection perception frameworks that require point or multi-view images, 3D-GS offers a source of high-quality training data without the need for extensive manual efforts. This aspect holds significant promise for future research endeavors.
In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters. Specifically, by leveraging the 2D foundational model SAM  [13], SA-GS can generate a segmented mask from a single input view based on the given point prompts. Starting from the obtained mask, we automatically generate multi-view masks and achieve consistent 3D segmentation via the proposed view-wise label assignment. In addition, to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme. Moreover, our segmented 3D Gaussians can be seamlessly employed in collision detection and scene editing tasks, like removal, translation, rotation, etc. In summary, the contributions of this paper are as follows.
  • We propose a simple yet effective method for object segmentation in 3D Gaussians without any training process and learnable parameters.
  • Efficient collision detection and scene editing, such as object removal, translation, and rotation within 3D scenes, can be performed readily based on the segmentation of SA-GS.
  • We conduct extensive segmentation experiments on a considerable amount of 3D scenes, demonstrating the effectiveness of our proposed method.
2 Related Work
2.1 3D Gaussian Splitting
3D Gaussian Splatting [11] is a technique for scene reconstruction, which has been shown to be an alternative 3D representation as NeRF, benefiting from both its high-quality rendering results and real-time rendering speed. Recent research on this technique contains applying 3D Gaussian in dynamic scene and combining it with diffusion model to achieve 3D generation. Dynamic 3D Gaussians [17] extends Gaussian Splatting to dynamic scenes via tracking 3D objects modeled as a set of 3D Gaussians. DreamGaussian [29], GaussianDreamer [32] and GSGEN [5] combine 3D Gaussians with diffusion models to generate high-quality 3D assets. In this work, we for the first time lift 2D foundation model-SAM[13] to 3D Gaussians in order to achieve 3D object segmentation in this Gaussian space.
2.2 Segmentation in NeRFs
Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks. Many researchers have explored how to segment objects in 3D using NeRFs, for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding. Some methods, such as Semantic-NeRF [33], NVOS [26], and SA3D [3], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks. Other methods, such as N3F [30], DFF [30], LERF [12], and ISRF [8], learn additional feature fields that are aligned with NeRFs, and use 2D visual features from pre-trained models or language embeddings to query the 3D features. These methods usually require modifying or retraining the original NeRF models or training other specific parameters to obtain the 3D segmentation. However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc. Benefiting from the purely explicit representation of 3D Gaussian, we can directly achieve 3D segmentation via an interactive way without any learned parameters and training process, and can apply for many other applications.
Figure 2: Pipeline of our proposed SA-GS. (a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
3 Method
In this section, we first present the preliminary in 3D Gaussian Splatting in Sec. 3.1 for clear understanding. We then define the problem and task in Sec. 3.2. Sec. 3.3, Sec. 3.4 and Sec. 3.5 elaborate on the technical details.
3.1 Preliminary: 3D Gaussian Splatting
3D Gaussian Splatting (3D GS) [11] is an emerging method for real-time radiance field rendering. It has been proven effective in Novel View Synthesis with high rendering quality as NeRF and real-time rendering speed. 3D GS represents scenes with a set of 3D Gaussians. Specifically, each 3D Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma$, an opacity value $\alpha$, and spherical harmonics (SH).
To render an image, it uses the splatting rendering pipeline, where 3D Gaussians are projected onto the 2D image plane. The projection transforms 3D Gaussians into 2D Gaussians in the image plane. All 2D Gaussians are blended together by the $\alpha$-blending algorithm to generate the color:
$\mathbf{c}=\sum_{i\in N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
During the $\alpha$-blending process, for each 2D Gaussian, only the 2D points with probability density larger than a certain threshold are calculated. This means a 2D Gaussian and 3D Gaussian can be intuitively regarded as a 2D ellipse and a 3D ellipsoid respectively. Empirically, for an axis of the ellipse, its length is set to $3\sigma$, where $\sigma$ is the square root the of variance in the axis.
3.2 Problem Definition
Given a set of trained 3D gaussians $\mathbb{G}=\{\mathbf{g}_{0},\mathbf{g}_{1},\cdots,\mathbf{g}_{n}\}$ and a random initial view $\mathbf{v}_{0}$, users could offer 2D point prompt set $\mathbb{P}_{2D}=\{\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{m}\}$ to specify a 2D object in view $\mathbf{v}_{0}$. Our algorithm is supposed to segment the corresponding 3D object $\mathbb{O}$ in $\mathbb{G}$ according to the human prompts, where $\mathbb{O}$ is a subset of $\mathbb{G}$.
Let $\mathbf{m}_{i}$ denote the projected binary mask of $\mathbb{O}$ in $i$-th view, an accurate segmentation $\mathbb{O}$ means $\mathbf{m}_{i}$ equals to $\mathbf{m}_{i}^{*}$ for $\forall i\in\{0,1,\cdots,n\}$. Here $\mathbf{m}_{i}^{*}$ is the ground truth mask of $\mathbb{O}$ in $i$-th view. Different from the conventional 2D segmentation mask in the image or 3D segmentation task in the point cloud, there is no ground truth for 3D Gaussians. Thus, our algorithm is designed to minimize the difference between $\mathbf{m}_{i}$ and $\mathbf{m}_{i}^{*}$.
3.3 Segment 3D Gaussians with 2D Mask
3D Prompts for Multiview Masks Generation
By the definition in Sec. 3.2, users are given the first rendered view to specify the target object. However, only the first view is far from sufficient to segment the target object in 3D space. So here we first generate multiview masks to aid the 3D segmentation. With multiple masks from different views, a 3D object can be segmented by the intersection of the corresponding frustum of these masks.
The core of obtaining masks is to generate 2D prompt points in each view. Denoting the $i$-th 2D prompt point in the first given view $\mathbf{v}_{0}$ as $\mathbf{p}_{i}^{0}$, we define the corresponding 3D prompt $\mathbf{p}_{i}^{3D}$ as:
$\operatorname*{arg\,min}_{\mu}\{d(\mu),d(\mu)>0\mid\mu\in\mathbb{G},\|\mathbf{P}_{0}\mu-\mathbf{p}_{i}^{0}\|_{1}<\epsilon\},$ (2)
where $d(\mu)$ is the depth of Gaussian center $\mu$ and $\mathbf{P}_{0}$ is the projection for initial view $\mathbf{v}_{0}$. Thus $\mathbf{P}_{0}\mu$ is the position of $\mu$ in view $\mathbf{v}_{0}$. Eq. 2 indicates that the corresponding 3D prompt of $\mathbf{p}_{i}$ is the center of a certain 3D Gaussian. This center meets two requirements: (1) it has a similar projected position with $\mathbf{p}_{i}$ with a Manhattan distance less than $\epsilon$, and (2) if there are multiple 3D Gaussian centers satisfying the first requirement, the one with the smallest positive depth is selected as the 3D prompt.
For all the 2D prompt points in the first view, we could get a set of 3D prompt by Eq. 2. Then for another view $\mathbf{v}_{i}$, we project these 3D prompts into the 2D plane, resulting in 2D prompts in the view $\mathbf{v}_{i}$. In this way, we obtain 2D prompts in all views, and all masks are obtained by prompting the SAM.
View-wise Label Assignment
With all the masks, we proceed to assign binary labels to each 3D Gaussian. In particular, we maintain a matrix $\mathbf{L}$, whose element $\mathbf{L}_{ij}$ is defined by
$\mathbf{L}_{ij}=\begin{cases}1&\text{if }\mathbf{P}_{j}\mu_{i}\in\mathbf{m}_{j},\\ 0&\text{if }\mathbf{P}_{j}\mu_{i}\notin\mathbf{m}_{j},\end{cases}$ (3)
where $\mu_{i}$ is the $i$-th Gausssian center of the scene and $\mathbf{m}_{j}$ is the foreground mask in $j$-th view. $\mathbf{P}_{j}$ stands for the projection matrix of $j$-th view.
Figure 3: Illustration of the Gaussian Decomposition process.
3.4 Gaussian Decomposition
Eq. 3 assign labels to 3D Gaussians only by the projected position of its center. However, 3D Gaussian has non-negligible spatial volume, so those Gaussians projected to the mask boundary usually have a part out of the boundary, greatly increasing the roughness of boundaries. A straightforward solution is directly removing the Gaussians across mask boundaries. However, such a solution greatly damages the 3D structures of the object.
To address this issue, we propose the Gaussian Decomposition to mitigate the boundary roughness while maintaining the 3D structures as complete as possible. Figure. 3 illustrates the basic idea of Gaussian decomposition. It has two basic steps to achieve the Gaussian decomposition: (1) For each 3D Gaussian, we obtain the corresponding 2D Gaussian by projection and mark it as a boundary Gaussian if one of its long-axis endpoints is outside of the 2D mask. Then, as shown in Figure. 3 we denote the two ends of the long axis of a 2D Gaussian as $A^{*}$ and $B^{*}$, and the intersection of $A^{*}B^{*}$ and mask boundary is $O^{*}$. $A$, $B$, and $O$ are the corresponding 3D points in the long axis of the 3D Gaussian. Assuming $A^{*}$ is in the mask while $B^{*}$ is out of the mask, we define
$\lambda_{3D}=\frac{OA}{AB},$ (4)
$\lambda_{2D}=\frac{O^{*}A^{*}}{A^{*}B^{*}}.$ (5)
(2) After obtaining the boundary Gaussians, we solve the ratio $\lambda_{3D}$ by first calculating $\lambda_{2D}$ and then decompose the original 3D Gaussian according to $\lambda_{3D}$.
However, the transformation from $\lambda_{2D}$ to $\lambda_{3D}$ is not straightforward because the perspective projection from 3D space to 2D space is not affine, which means the ratios are different. Fortunately, 3D Gaussian Splatting leverages local affine approximation to simplify the rendering process. It projects 3D Gaussian to the 2D plane by
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$ (6)
where $\Sigma$, $\Sigma^{\prime}$ are the covariance matrix of the 3D and 2D gaussian distribution, respectively. $W$ is the projection transformation and $J$ is the Jocobian of affine approximation derived in EWA algorithm [34]. Eq. 6 indicates that Gaussian Splatting simplifies the perspective projection to an affine projection, thus decomposition scaling ratio $\lambda_{3D}$ in 3D space is equivalent to the ratio $\lambda_{2D}$ in the 2D plane.
Let $\mathbf{g}$ denote a 3D Gaussian across the boundary. Its scale in long axis and the 3D center are defined as $s$ and $\mathbf{\mu}$, respectively. We have
$s^{\prime}=\lambda_{2D}s,$ (7)
$\mathbf{\mu}^{\prime}=\mathbf{\mu}+\frac{1}{2}\left(s-\lambda_{2D}s\right)\mathbf{e},$ (8)
where $\mathbf{e}$ is the unit vector pointing from the 3D Gaussian center to the in-mask endpoint of the long axis. The decomposed Gaussian $\mathbf{g}^{\prime}$ adopt $\mathbf{\mu}^{\prime}$ and $s^{\prime}$ as the new center and long-axis scale, maintaining other properties unchanged. Another decomposed Gaussian outside of the mask is removed.
3.5 Multiview Label Voting
So far, every 3D Gaussian including the decomposed one has a list of binary labels $\mathbf{L}_{i}$, using the label assignment in Sec. 3.3. Leveraging the assigned labels, here we adopt a simple yet effective heuristic rule to determine if a 3D Gaussian $\mathbf{g}_{i}$ belongs to the target 3D object. In particular, we first define the confidence score $s_{i}$ of $g_{i}$ as
$s_{i}=\frac{1}{N}\sum_{j=0}^{N-1}L_{ij},$ (9)
where $N$ is the number of views. Then we adopt a threshold $\tau$ and regard a Gaussian with a score higher than $\tau$ as positive.
4 Experimets
Figure 4: Visualization of segmentation results in different scenes (LERF-figurines [12], LERF-dozer-nerfgun-waldo [12], SPIn-NeRF-Orchids [21]). We enlarge the boxed area on the right for a better visualization.
4.1 Datasets
We choose different datasets to testify our method, including LLFF [19], Mip-NeRF 360 [2], LERF [12], and some test scenes from the 3D Gaussian Splatting [11]. These datasets contain both small indoor objects and large outdoor scenes, which are very complex and challenging. For quantitative experiments, because there is no existing benchmark that can be used in 3D Gaussian space, we use the SPIn-NeRF [21] dataset with 2D ground-truth for evaluation.
4.2 Quantitative results
Scenes Single View MVSeg SA3D Ours
IoU Acc IoU Acc IoU Acc IoU Acc
Orchids 79.4 96.0 92.7 98.8 83.6 96.9 85.4 97.5
Ferns 95.2 99.3 94.3 99.2 97.1 99.6 92.0 98.9
Room 73.4 96.5 95.6 99.4 88.2 98.3 86.5 98.1
Horns 85.3 97.1 92.8 98.7 94.5 99.0 91.1 98.4
Fortress 94.1 99.1 97.7 99.7 98.3 99.8 96.6 99.5
Fork 69.4 98.5 87.9 99.5 89.4 99.6 83.4 99.3
Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 92.0 99.0
Truck 37.9 77.9 85.2 95.1 90.8 96.7 93.0 97.9
Lego 76.0 99.1 74.9 99.2 92.2 99.8 88.4 99.7
Mean 74.1 95.2 90.5 98.8 91.9 98.8 89.9 98.7
Table 1: Quantitative results on SPIn-NeRF dataset
We first conduct experiments on the SPIn-NeRF [21] dataset for quantitative analysis. Given a set of images of the scene, we follow the process described in Section 3 to obtain the segmentation of the target object in 3D Gaussian Space. The segmented 3D Gaussians are used to render 2D masks in other views. Finally, we calculate the IoU and Accuracy between these rendered and the ground-truth masks. Results can be seen in Table 1.
Figure 5: Visualization examples of scene editing after the object segmentation. We offer three scene editing examples: removal, translation, and rotation.
In the comparison, it is noteworthy that both the MVSeg [21] and the SA3D [3] require additional parameters and a computation-costly training process. By contrast, the “Single view” [3] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process. In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters. Results show that our approach can achieve comparable segmentation quality to the learnable MVSeg and SA3D methods. In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg. When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.8% IoU and +3.5% Acc. These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
Scene room truck fortress pinecone
Text “the table” “the truck” “the fortress” “the pinecone”
IoU 86.3 93.7 92.8 86.5
Acc 97.9 97.8 98.8 97.0
Table 2: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
Next, we replace the above point prompts with text prompts corresponding to the objects. Given a text input, we prompt the Grounding DINO [16] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks. Then, following the segmentation process in Section 3, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input. Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks. Quantitative results are shown in Table 2. For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts. These results demonstrate our method has the ability to combine multi-modal prompts as input.
4.3 Qualitative results
We conduct four kinds of tasks to demonstrate the potential application and qualitative performance of our approach, including point-guided segmentation, text-guided segmentation, scene editing, and collision detection.
Point-Guided Segmentation:
We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image. The results are shown in Figure 4. Our approach demonstrates stable performance in diverse scenes and target objects, including multiple instance segmentation (1st row), salient object segmentation (2nd row), and partial segmentation (3rd row). Even when the target objects are of extremely small scales (the first row), our approach can still produce high-quality segmentation.
Scene w/ GD w/o GD Delete
IoU Acc IoU Acc IoU Acc
Orchids 85.4 97.5 82.2 96.8 78.2 95.4
Ferns 92.0 98.9 89.2 98.4 89.8 98.5
Room 86.5 98.1 81.3 97.2 85.4 97.9
Horns 91.1 98.4 83.2 96.5 84.8 97.6
Fortress 96.5 99.4 88.5 98.1 82.3 96.7
Fork 83.4 99.3 81.8 99.2 79.9 99.1
Pinecone 92.1 98.9 91.6 98.9 82.9 97.5
Truck 94.0 97.9 93.4 97.8 91.4 96.8
Lego 88.4 99.7 88.4 99.6 82.9 99.4
mean 89.9 98.7 86.6 98.1 84.2 97.7
Table 3: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset. The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD). The last column represents the results of directly removing the gaussians across mask boundaries.
Text-Guided Segmentation:
Next, we further conduct 3D object segmentation experiments guided by text prompts. We follow the same process as in Table 2 experiments to conduct visualization experiments. In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”. Results in Figure 6 show that our approach can achieve accurate segmentation results by simply providing object names, which demonstrates the potential of our method in combining with language models. Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
Figure 6: 3D segmentation results with text prompts in Mip-NeRF360-garden [2].
Scene Editing:
Scene editing is a basic application for 3D reconstructed scenes. However, it’s quite difficult to do this without being able to locate the specific objects. This task demonstrates the ability of our approach to help edit the 3D scenes. Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them. Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Figure 5. The instances in 1st column with red bounding boxes are objects to be segmented. It can be seen that with objects segmented in 3D Gaussians, they can be translated and rotated in any direction in the scene. After the removal of segmented objects, original scenes can still keep intact.
Collision Detection:
Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators. In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space. We choose two scenes from SPIn-NeRF dataset. Following the process of our segmentation method, we can obtain the corresponding segmented 3D Gaussian points in the scene. To this end, we use the Quickhull [1] algorithm to build the collision mesh upon our segmented objects. The results in Figure 7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
Number of views 5(10%) 10(20%) 21(50%) 42(100%)
IoU on Fortress 91.16 92.11 93.82 96.55
Number of views 25(10%) 50(20%) 125(50%) 251(100%)
IoU on truck 90.27 90.97 92.11 93.48
Table 4: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
Figure 7: Collision mesh for segmented 3D objects using QuickHull algorithm.
$\tau$ 0.3 0.5 0.6 0.7 0.9
IoU 61.7 82.7 93.4 90.6 53.9
Table 5: Ablation on confidence score threshold $\tau$ on the “truck” scene.
Figure 8: The influence of different view numbers on LERF-fern.
4.4 Ablation Study
Figure 9: The influence of different view directions.
Selection of Views:
In this section, we study the influence of selected 2D views on 3D Gaussian segmentation. Figure 8 demonstrates the segmentation quality using different numbers of views. It is noteworthy that the leaves in the “fern” scene are very tiny and challenging. Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views. Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase. It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach. In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
Figure 9 studies the influence of view directions on final segmentation results. We choose the truck with 360-degree views. The upper row shows the result of using roughly a single direction. In this case, the background is mistakenly incorporated into the segmentation. By contrast, with the same number of views but variant directions, the approach achieves high-quality clean segmentation results. This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
Gaussian Decomposition:
Gaussian Decomposition is proposed to address the issue of roughness boundaries of 3D segmented objects, which results from the non-negligible spatial sizes of 3D Gaussian located at the boundary. We conduct both quantitative and qualitative experiments to verify the effectiveness of our approach. For quantitative ablation experiments, Table 3 shows the results. We compare our proposed Gaussian Decomposition scheme with two other processing methods, one for segmentation without any special handling (the 3rd column), and the other for directly removing such gaussians (the last column). The experiments are conducted on SPIn-NeRF dataset, following the same evaluation process as Table 1. Compared with the others, our proposed approach outperforms them on all scenes with +3.3% and +5.7% IoU, respectively, demonstrating the effectiveness of Gaussian Decomposition. Besides, it can seen that directly removing these gaussians cannot decrease the roughness of mask boundaries, leading to even 2.4% lower IoU values.
Figure 10 shows the visualization results for comparison. We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column). This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy. Though this idea is simple, the improvement is obvious.
Figure 10: Ablation results on Gaussian Decomposition (GD) on the LERF-pinecone.
Hyper-Parameters:
In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Section 3.5. As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects. For scenes with some views that may cover only part of the objects to be segmented (i.e. half of the truck body in the “truck” scene), a small $\tau$ value is preferred. In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
5 Conclusion
In this paper, we present a novel approach for object segmentation in 3D Gaussians, which performs an interactive procedure without any training process and learnable parameters. Based on a trained 3D Gaussian representation and clicked points on first view, we propose the multiview mask generation and view-wise label assignment scheme to generalize SAM to achieve 3D segmentation. In addition, we provide the Gaussian Decomposition to mitigate the boundary roughness issue of segmented masks, and the cross-view label voting approach to assign labels from different views. Extensive segmentation experiments show that our method can effectively obtain high-quality object segmentation results in 3D Gaussians, which can be easily applied for scene editing and collision detection tasks. Overall, we hope our work can inspire more future work in the area of 3D Gaussian representation.

### v3

@@
-Segment Anything in 3D Gaussians
+
 Abstract
-3D Gaussian Splatting has emerged as an alternative 3D representation of Neural Radiance Fields (NeRFs), benefiting from its high-quality rendering results and real-time rendering speed.
-Considering the 3D Gaussian representation remains unparsed, it is necessary first to execute object segmentation within this domain.
-Subsequently, scene editing and collision detection can be performed, proving vital to a multitude of applications, such as virtual reality (VR), augmented reality (AR), game/movie production, etc.
-In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters.
-We refer to the proposed method as SA-GS, for Segment Anything in 3D Gaussians.
-Given a set of clicked points in a single input view, SA-GS can generalize SAM to achieve 3D consistent segmentation via the proposed multi-view mask generation and view-wise label assignment methods.
-We also propose a cross-view label-voting approach to assign labels from different views.
-In addition, in order to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme.
-Extensive experiments demonstrate that SA-GS achieves high-quality 3D segmentation results, which can also be easily applied for scene editing and collision detection tasks.
-Codes will be released soon.
+3D Gaussian Splatting has emerged as an alternative 3D representation for novel view synthesis, benefiting from its high-quality rendering results and real-time rendering speed.
+However, the 3D Gaussians learned by 3D-GS have ambiguous structures without any geometry constraints.
+This inherent issue in 3D-GS leads to a rough boundary when segmenting individual objects.
+To remedy these problems, we propose SAGD, a conceptually simple yet effective boundary-enhanced segmentation pipeline for 3D-GS to improve segmentation accuracy while preserving segmentation speed.
+Specifically, we introduce a Gaussian Decomposition scheme, which ingeniously utilizes the special structure of 3D Gaussian, finds out, and then decomposes the boundary Gaussians.
+Moreover, to achieve fast interactive 3D segmentation, we introduce a novel training-free pipeline by lifting a 2D foundation model to 3D-GS.
+Extensive experiments demonstrate that our approach achieves high-quality 3D segmentation without rough boundary issues, which can be easily applied to other scene editing tasks.
+Our code is public available at https://github.com/XuHu0529/SAGS.
+Keywords:
+3D Gaussian Splatting 3D Segmentation Boundary Issues
 1 Introduction
-Figure 1: We propose an effective 3D object segmentation method in 3D Gaussian without any training process.
-The segmented mask contributes to the realization of various applications, including collision detection, translation, rotation and removal, etc.
 3D scene understanding is a challenging and crucial task in computer vision and computer graphics, which involves scene reconstruction from images or videos and the perception of a given 3D real-world environment.
@@
 Although current methods have attained noteworthy success in 3D scene understanding, the time-consuming nature of NeRF and the high costs associated with 3D data collection pose challenges for scaling up these approaches.
-Alternatively, SA3D  [3] addresses the aforementioned challenges by enhancing the 2D foundational model (i.e.
-SAM  [13]) with 3D perception through the incorporation of a 3D representation model.
-SA3D demonstrates notable success by leveraging NeRF to generate 2D views and aligning 2D segmentation results with the 3D scene.
-However, the substantial training and rendering time costs associated with NeRF make it impractical for large-scale 3D scenes.
-Consequently, in this work, we focus on 3D Gaussian Splatting (3D-GS) [11] as a prospective method for modeling static 3D scenes.
+Figure 1: (a)The training of 3D-GS doesn’t consider the structure of objects, leading to the ambiguous geometry; (b)Direct segmentation without Gaussian Decomposition processing will result in rough boundary segmentation; (3) The recent SAGA also has incomplete segmentation caused by the same issue; (d) Our full pipeline considers this issue and can achieve better segmentation.
+Recently, 3D Gaussian Splatting (3D-GS) [12] is emerging as a prospective method for modeling static 3D scenes.
 3D-GS characterizes intricate scenes by employing numerous colored 3D Gaussians, rendering them into camera views through splatting-based rasterization.
 Through differentiable rendering and gradient-based optimization, the positions, sizes, rotations, colors, and opacities of these Gaussians can be finely tuned to accurately represent the 3D scene, availing for the comprehension of a 3D environment.
-Benefiting from the properties of high-quality and real-time rendering in 3D-GS, it is impressive to use these explicit representations to capture the shape and appearance of complex objects.
-Specifically, based on the 3D Gaussian volume, we can effectively achieve scene perception via equipping the 2D foundation model as SA3D  [3] statement.
-Different from SA3D leveraging implicit NeRF as a 3D prior to connecting multiple 2D views, our method (SA-GS) applies effect and explicit 3D-GS representation to perform 3D object segmentation.
-Subsequently, scene editing and collision detection can be executed readily, as shown in Figure 1, which is essential to various applications, such as virtual reality(VR), augmented reality(AR), and games/movie production.
-Moreover, in contrast to the existing labor-intensive data collection perception frameworks that require point or multi-view images, 3D-GS offers a source of high-quality training data without the need for extensive manual efforts.
-This aspect holds significant promise for future research endeavors.
-In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters.
-Specifically, by leveraging the 2D foundational model SAM  [13], SA-GS can generate a segmented mask from a single input view based on the given point prompts.
-Starting from the obtained mask, we automatically generate multi-view masks and achieve consistent 3D segmentation via the proposed view-wise label assignment.
-In addition, to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme.
-Moreover, our segmented 3D Gaussians can be seamlessly employed in collision detection and scene editing tasks, like removal, translation, rotation, etc.
-In summary, the contributions of this paper are as follows.
-  • We propose a simple yet effective method for object segmentation in 3D Gaussians without any training process and learnable parameters.
-  • Efficient collision detection and scene editing, such as object removal, translation, and rotation within 3D scenes, can be performed readily based on the segmentation of SA-GS.
-  • We conduct extensive segmentation experiments on a considerable amount of 3D scenes, demonstrating the effectiveness of our proposed method.
+Segmentation in 3D-GS is quite a natural pathway to scene understanding for its explicit representation.
+Recent works [3, 34, 37, 16] have been proposed to achieve segmentation in 3D-GS via lifting 2D foundation models in a learnable way.
+They share the same ideas to distill other features or identity encodings generated from 2D foundation models to 3D Gaussian fields.
+However, the Gaussians learned by 3D-GS are ambiguous without any geometry constraint, leading to the issue of boundary roughness.
+A single Gaussian might correspond to multiple objects, complicating the task of accurately segmenting individual objects (shown in Fig.
+1 (a)).
+As shown in Fig.
+1 (b), direct segmentation will leave rough edges at the boundary.
+This is because these Gaussians across multiple objects will be left if no additional processing is performed.
+In addition, even if SAGA [3] uses filtering and growing post-processing for segmented 3D Gaussians, its feature matching method also has the limitation of incomplete boundary, as shown in Fig.
+1 (c).
+We present our boundary-enhanced segmentation method, an interactive training-free pipeline for efficient and effective segmentation in 3D-GS without rough boundary issues.
+By leveraging the 2D foundation model SAM [14], our method can generate a segmented mask from a single input view based on the given prompts.
+Starting from the obtained mask, our method automatically generates multi-view masks and achieves consistent 3D segmentation via the proposed assignment strategy.
+Specifically, to resolve the inherent boundary issues, we incorporate a simple but effective Gaussian Decomposition scheme, ingeniously utilizing the special structure of 3D Gaussian.
+Since the boundary roughness issue results from the non-negligible spatial sizes of 3D Gaussians located at the boundary, our key insight is to find and then decompose the original boundary Gaussians according to our proposed principles.
+Consequently, our method eliminates the inherent issues and achieves more complete segmentation quickly and efficiently, as shown in Fig.
+1 (d).
+In summary, the contributions of this paper are as follows:
+  • We propose a simple yet effective training-free pipeline for segmentation in 3D Gaussians without any learnable parameters;
+  • We incorporate the Gaussian Decomposition module to mitigate the boundary roughness issues in 3D segmentation resulting from inherent 3D-GS geometry;
+  • Extensive segmentation experiments on a considerable amount of 3D scenes and editing applications demonstrate the effectiveness of our proposed method.
 2 Related Work
@@
 DreamGaussian [31], GaussianDreamer [35] and GSGEN [6] combine 3D Gaussians with diffusion models to generate high-quality 3D assets.
-In this work, we for the first time lift 2D foundation model-SAM[13] to 3D Gaussians in order to achieve 3D object segmentation in this Gaussian space.
-2.2 Segmentation in NeRFs
+2.2 Segmentation in Radiance Fields
 Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks.
@@
 However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc.
-Benefiting from the purely explicit representation of 3D Gaussian, we can directly achieve 3D segmentation via an interactive way without any learned parameters and training process, and can apply for many other applications.
-Figure 2: Pipeline of our proposed SA-GS.
+Inspired by the real-time 3D-GS, recent works [3, 34, 37, 16] have been proposed to achieve segmentation in 3D-GS in a learnable way.
+SAGA [3] and Featrue 3DGS [37] distill the knowledge embedded in the SAM decoder into the feature field of 3D-GS, and Gaussian-Grouping [34] supervises the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by SAM.
+However, no research has been proposed to solve the rough boundary issues in segmentation resulting from the inherent properties of 3D Gaussians.
+In this work, we are the first to propose a novel Gaussian Decomposition scheme to solve this issue, incorporated into our effective training-free segmentation pipeline.
+Figure 2: Pipeline of our proposed method.
 (a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
@@
 In this section, we first present the preliminary in 3D Gaussian Splatting in Sec.
-3.1 for clear understanding.
-We then define the problem and task in Sec.
+3.1 for clear understanding and then define the problem and task in Sec.
 3.2.
-Sec.
-3.3, Sec.
+Our basic training-free pipeline consists of Sec.
 3.3 and Sec.
-3.5 elaborate on the technical details.
+3.5.
+The details of Gaussian Decomposition are described in Sec.
+3.4.
+The pipeline of our method can be seen in Fig.
+2.
 3.1 Preliminary: 3D Gaussian Splatting
@@
 3.3 Segment 3D Gaussians with 2D Mask
-3D Prompts for Multiview Masks Generation
+3D Prompts for Multiview Masks Generation.
 By the definition in Sec.
@@
 In this way, we obtain 2D prompts in all views, and all masks are obtained by prompting the SAM.
-View-wise Label Assignment
-With all the masks, we proceed to assign binary labels to each 3D Gaussian. In particular, we maintain a matrix $\mathbf{L}$, whose element $\mathbf{L}_{ij}$ is defined by
+View-wise Label Assignment. With all the masks, we proceed to assign binary labels to each 3D Gaussian. In particular, we maintain a matrix $\mathbf{L}$, whose element $\mathbf{L}_{ij}$ is defined by
 $\mathbf{L}_{ij}=\begin{cases}1&\text{if }\mathbf{P}_{j}\mu_{i}\in\mathbf{m}_{j},\\ 0&\text{if }\mathbf{P}_{j}\mu_{i}\notin\mathbf{m}_{j},\end{cases}$ (3)
@@
 3.4 Gaussian Decomposition
-Eq.
-3 assign labels to 3D Gaussians only by the projected position of its center.
+After obtaining the prompt points of each view, we could assign labels to 3D Gaussians by the projected position of its center.
 However, 3D Gaussian has non-negligible spatial volume, so those Gaussians projected to the mask boundary usually have a part out of the boundary, greatly increasing the roughness of boundaries.
@@
 However, such a solution greatly damages the 3D structures of the object.
-To address this issue, we propose the Gaussian Decomposition to mitigate the boundary roughness while maintaining the 3D structures as complete as possible.
-Figure.
-3 illustrates the basic idea of Gaussian decomposition.
-It has two basic steps to achieve the Gaussian decomposition: (1) For each 3D Gaussian, we obtain the corresponding 2D Gaussian by projection and mark it as a boundary Gaussian if one of its long-axis endpoints is outside of the 2D mask.
-Then, as shown in Figure.
-3 we denote the two ends of the long axis of a 2D Gaussian as $A^{*}$ and $B^{*}$, and the intersection of $A^{*}B^{*}$ and mask boundary is $O^{*}$.
-$A$, $B$, and $O$ are the corresponding 3D points in the long axis of the 3D Gaussian.
-Assuming $A^{*}$ is in the mask while $B^{*}$ is out of the mask, we define
+[…]
 $\lambda_{3D}=\frac{OA}{AB},$ (4)
@@
 6 indicates that Gaussian Splatting simplifies the perspective projection to an affine projection, thus decomposition scaling ratio $\lambda_{3D}$ in 3D space is equivalent to the ratio $\lambda_{2D}$ in the 2D plane.
-Let $\mathbf{g}$ denote a 3D Gaussian across the boundary. Its scale in long axis and the 3D center are defined as $s$ and $\mathbf{\mu}$, respectively. We have
+[…]
 $s^{\prime}=\lambda_{2D}s,$ (7)
@@
 4 Experimets
-Figure 4: Visualization of segmentation results in different scenes (LERF-figurines [12], LERF-dozer-nerfgun-waldo [12], SPIn-NeRF-Orchids [21]).
-We enlarge the boxed area on the right for a better visualization.
 4.1 Datasets
@@
 4.2 Quantitative results
-Scenes Single View MVSeg SA3D Ours
-IoU Acc IoU Acc IoU Acc IoU Acc
-Orchids 79.4 96.0 92.7 98.8 83.6 96.9 85.4 97.5
-Ferns 95.2 99.3 94.3 99.2 97.1 99.6 92.0 98.9
-Room 73.4 96.5 95.6 99.4 88.2 98.3 86.5 98.1
-Horns 85.3 97.1 92.8 98.7 94.5 99.0 91.1 98.4
-Fortress 94.1 99.1 97.7 99.7 98.3 99.8 96.6 99.5
-Fork 69.4 98.5 87.9 99.5 89.4 99.6 83.4 99.3
-Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 92.0 99.0
-Truck 37.9 77.9 85.2 95.1 90.8 96.7 93.0 97.9
-Lego 76.0 99.1 74.9 99.2 92.2 99.8 88.4 99.7
-Mean 74.1 95.2 90.5 98.8 91.9 98.8 89.9 98.7
-Table 1: Quantitative results on SPIn-NeRF dataset
+Table 1: Quantitative results on SPIn-NeRF dataset.
+“Singe view” denotes projecting the 2D segmentation result to 3D simply, thus we don’t consider its time cost.
+There is extra time cost (about 10mins) for SAGA to distill 2D knowledge to 3D-GS, while other methods don’t need this process.
+Scenes Single View [4] MVSeg [23] SA3D [4] SAGA [3] Ours
+IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc
+Orchids 79.4 96.0 92.7 98.8 83.6 96.9 - - 85.4 97.5
+Ferns 95.2 99.3 94.3 99.2 97.1 99.6 - - 92.0 98.9
+Room 73.4 96.5 95.6 99.4 88.2 98.3 - - 86.5 98.1
+Horns 85.3 97.1 92.8 98.7 94.5 99.0 - - 91.1 98.4
+Fortress 94.1 99.1 97.7 99.7 98.3 99.8 - - 96.6 99.5
+Fork 69.4 98.5 87.9 99.5 89.4 99.6 - - 83.4 99.3
+Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 - - 92.0 99.0
+Truck 37.9 77.9 85.2 95.1 90.8 96.7 - - 93.0 97.9
+Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 88.4 99.7
+Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 89.9 98.7
+Time cost - 3-6mins 2-10mins 0.08s-0.9s within 15s
+(/object) (+10mins)
 We first conduct experiments on the SPIn-NeRF [23] dataset for quantitative analysis.
@@
 Results can be seen in Table 1.
-Figure 5: Visualization examples of scene editing after the object segmentation. We offer three scene editing examples: removal, translation, and rotation.
 In the comparison, it is noteworthy that both the MVSeg [23] and the SA3D [4] require additional parameters and a computation-costly training process.
+While SAGA [3] shows nearly one-thousandth of the time compared with them, it needs extra training time (10 minutes per scene) to distill the knowledge embedded in the SAM decoder into the feature field.
 By contrast, the “Single view” [4] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process.
 In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters.
-Results show that our approach can achieve comparable segmentation quality to the learnable MVSeg and SA3D methods.
+Results show that our approach can achieve comparable segmentation quality to the MVSeg and SA3D methods and better results than SAGA.
+The time cost of ours is much less than MVSeg and SA3D.
+Considering the training time of SAGA, the average segmentation cost of a single object is basically the same as our method.
 In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg.
@@
 These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
+Table 2: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
 Scene room truck fortress pinecone
@@
 Acc 97.9 97.8 98.8 97.0
-Table 2: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
+Figure 4: Qualitative results compared with SA3D [4] and SAGA [3] in different scenes (LERF-figurines [13], SPIn-NeRF-Orchids [23], LERF-dozer-nerfgun-waldo [13]).
+We enlarge the boxed area on the right for a better visualization.
 Next, we replace the above point prompts with text prompts corresponding to the objects.
@@
 For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts.
-These results demonstrate our method has the ability to combine multi-modal prompts as input.
+These results demonstrate our method can combine multi-modal prompts as input.
 4.3 Qualitative results
 We conduct four kinds of tasks to demonstrate the potential application and qualitative performance of our approach, including point-guided segmentation, text-guided segmentation, scene editing, and collision detection.
-Point-Guided Segmentation:
-We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image.
-The results are shown in Figure 4.
-Our approach demonstrates stable performance in diverse scenes and target objects, including multiple instance segmentation (1st row), salient object segmentation (2nd row), and partial segmentation (3rd row).
-Even when the target objects are of extremely small scales (the first row), our approach can still produce high-quality segmentation.
+Figure 5: 3D segmentation results with text prompts in Mip-NeRF360-garden [2].
+Point-Guided Segmentation: We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image.
+The results are shown in Fig.
+4.
+The first row shows the segmentation results of SA3D, SAGA and ours on LERF-figurines scene.
+The second row compares with SAGA, which distills the knowledge embedded in the SAM decoder into the feature field.
+Though the boundary issues in 3D-GS can be alleviated, incomplete segmentation occurs even after post-processing.
+In contrast, ours can solve the boundary problems while achieving more complete segmentation.
+More segmentation results are shown in the third row.
+Text-Guided Segmentation: We further conduct 3D object segmentation experiments guided by text prompts.
+We follow the same process as in Table 2 experiments to conduct visualization experiments.
+In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”.
+Results in Fig.
+5 show that our approach can achieve accurate segmentation results by simply providing object names, demonstrating our method’s potential in combining with language models.
+Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
+Comparison between 2D and 3D Segmentation: Our method improves the segmentation quality of SAM.
+Due to the sensitivity of SAM to the scene viewpoint, the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig.
+6).
+Table 3: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset.
+The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD).
+The last column represents the results of directly removing the gaussians across mask boundaries.
 Scene w/ GD w/o GD Delete
@@
 mean 89.9 98.7 86.6 98.1 84.2 97.7
-Table 3: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset.
-The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD).
-The last column represents the results of directly removing the gaussians across mask boundaries.
-Text-Guided Segmentation:
-Next, we further conduct 3D object segmentation experiments guided by text prompts.
-We follow the same process as in Table 2 experiments to conduct visualization experiments.
-In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”.
-Results in Figure 6 show that our approach can achieve accurate segmentation results by simply providing object names, which demonstrates the potential of our method in combining with language models.
-Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
-Figure 6: 3D segmentation results with text prompts in Mip-NeRF360-garden [2].
-Scene Editing:
-Scene editing is a basic application for 3D reconstructed scenes.
+Scene Editing: Scene editing is a basic application for 3D reconstructed scenes.
 However, it’s quite difficult to do this without being able to locate the specific objects.
@@
 Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them.
-Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Figure 5.
+Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Fig.
+7.
 The instances in 1st column with red bounding boxes are objects to be segmented.
@@
 After the removal of segmented objects, original scenes can still keep intact.
-Collision Detection:
-Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators.
+Figure 6: Comparison between 2D and 3D Segmentation.
+Collision Detection: Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators.
 In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space.
@@
 To this end, we use the Quickhull [1] algorithm to build the collision mesh upon our segmented objects.
-The results in Figure 7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
-Number of views 5(10%) 10(20%) 21(50%) 42(100%)
-IoU on Fortress 91.16 92.11 93.82 96.55
-Number of views 25(10%) 50(20%) 125(50%) 251(100%)
-IoU on truck 90.27 90.97 92.11 93.48
-Table 4: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
-Figure 7: Collision mesh for segmented 3D objects using QuickHull algorithm.
-$\tau$ 0.3 0.5 0.6 0.7 0.9
-IoU 61.7 82.7 93.4 90.6 53.9
-Table 5: Ablation on confidence score threshold $\tau$ on the “truck” scene.
-Figure 8: The influence of different view numbers on LERF-fern.
+The results in the last column of Fig.
+7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
+Figure 7: Visualization examples of scene editing after the object segmentation.
+We offer three scene editing examples: removal, translation, and rotation.
+We further provide collision mesh computed after segmentation, as shown in the last column.
 4.4 Ablation Study
-Figure 9: The influence of different view directions.
-Selection of Views:
-In this section, we study the influence of selected 2D views on 3D Gaussian segmentation.
-Figure 8 demonstrates the segmentation quality using different numbers of views.
-It is noteworthy that the leaves in the “fern” scene are very tiny and challenging.
-Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views.
-Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase.
-It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach.
-In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
-Figure 9 studies the influence of view directions on final segmentation results.
-We choose the truck with 360-degree views.
-The upper row shows the result of using roughly a single direction.
-In this case, the background is mistakenly incorporated into the segmentation.
-By contrast, with the same number of views but variant directions, the approach achieves high-quality clean segmentation results.
-This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
 Gaussian Decomposition:
@@
 Besides, it can seen that directly removing these gaussians cannot decrease the roughness of mask boundaries, leading to even 2.4% lower IoU values.
-Figure 10 shows the visualization results for comparison.
+Fig.
+8 shows the visualization results for comparison.
 We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column).
@@
 Though this idea is simple, the improvement is obvious.
+Table 4: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
+Number of views 5(10%) 10(20%) 21(50%) 42(100%)
+IoU on Fortress 91.16 92.11 93.82 96.55
+Number of views 25(10%) 50(20%) 125(50%) 251(100%)
+IoU on truck 90.27 90.97 92.11 93.48
+Table 5: Ablation on confidence score threshold $\tau$ on the “truck” scene.
+$\tau$ 0.3 0.5 0.6 0.7 0.9
+IoU 61.7 82.7 93.4 90.6 53.9
+Selection of Views:
+In this section, we study the influence of selected 2D views on 3D Gaussian segmentation.
+Fig.
+9 (b) demonstrates the segmentation quality using different numbers of views.
+It is noteworthy that the leaves in the “fern” scene are very tiny and challenging.
+Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views.
+Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase.
+It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach.
+In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
 Figure 8: Ablation results on Gaussian Decomposition (GD) on the LERF-pinecone.
+Fig.
+9 (a) study the influence of view directions on final segmentation results.
+We choose the truck with 360-degree views.
+The upper row shows the result of using roughly a single direction.
+In this case, the background is mistakenly incorporated into the segmentation.
+By contrast, with the same number of views but variant directions, the approach achieves high-quality clean segmentation results.
+This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
 Hyper-Parameters:
-In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Section 3.5.
+In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec.
+3.5.
 As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects.
@@
 In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
+5 Discussions and Limitations
+Through experiments, we found that our method does not perform well in objects where 3D Gaussians are very sparse, such as the LLFF-room [22] scene.
+The Gaussians of the table are notably sparse; even worse, the Gaussians representing the table surface are across different objects.
+Though our Gaussian Decomposition can somewhat remedy the boundary issue, it still suffers from extremely sparse points and leaves holes in 3D segmentation.
+We believe this limitation can be alleviated by future research in a more structured 3D-GS representation, yielding more accurate results.
+Figure 9: Ablation on view selection.
 6 Conclusion
-In this paper, we present a novel approach for object segmentation in 3D Gaussians, which performs an interactive procedure without any training process and learnable parameters.
-Based on a trained 3D Gaussian representation and clicked points on first view, we propose the multiview mask generation and view-wise label assignment scheme to generalize SAM to achieve 3D segmentation.
-In addition, we provide the Gaussian Decomposition to mitigate the boundary roughness issue of segmented masks, and the cross-view label voting approach to assign labels from different views.
-Extensive segmentation experiments show that our method can effectively obtain high-quality object segmentation results in 3D Gaussians, which can be easily applied for scene editing and collision detection tasks.
+We address the issue that rough/incomplete boundaries in segmenting 3D-GS and propose a novel Boundary-enhanced segmentation pipeline.
+Given input prompts on the first rendering view, our approach automatically generates multi-view masks and achieves consistent 3D segmentation via the proposed assignment strategy.
+Our Gaussian Decomposition module can effectively mitigate the boundary roughness issue of segmented objects resulting from the inherent geometry structure in 3D-GS.
+Extensive segmentation experiments show that our method can effectively obtain high-quality 3D object segmentation without boundary issues, and different scene-editing tasks demonstrate that our method can be easily applied to downstream applications.
 Overall, we hope our work can inspire more future work in the area of 3D Gaussian representation.

### v4

@@
-
+SAGD: Boundary-Enhanced Segment Anything in 3D Gaussian via Gaussian Decomposition
 Abstract
@@
 Extensive experiments demonstrate that our approach achieves high-quality 3D segmentation without rough boundary issues, which can be easily applied to other scene editing tasks.
-Our code is public available at https://github.com/XuHu0529/SAGS.
-Keywords:
-3D Gaussian Splatting 3D Segmentation Boundary Issues
-1 Introduction
+Our code is publicly available at https://github.com/XuHu0529/SAGS.
+Index Terms:
+3D gaussian splatting, 3D segmentation, boundary issues.
+I Introduction
 3D scene understanding is a challenging and crucial task in computer vision and computer graphics, which involves scene reconstruction from images or videos and the perception of a given 3D real-world environment.
@@
   • Extensive segmentation experiments on a considerable amount of 3D scenes and editing applications demonstrate the effectiveness of our proposed method.
-2 Related Work
-2.1 3D Gaussian Splitting
-3D Gaussian Splatting [12] is a technique for scene reconstruction, which has been shown to be an alternative 3D representation as NeRF, benefiting from both its high-quality rendering results and real-time rendering speed.
-Recent research on this technique contains applying 3D Gaussian in dynamic scene and combining it with diffusion model to achieve 3D generation.
-Dynamic 3D Gaussians [19] extends Gaussian Splatting to dynamic scenes via tracking 3D objects modeled as a set of 3D Gaussians.
-DreamGaussian [31], GaussianDreamer [35] and GSGEN [6] combine 3D Gaussians with diffusion models to generate high-quality 3D assets.
-2.2 Segmentation in Radiance Fields
-Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks.
-Many researchers have explored how to segment objects in 3D using NeRFs, for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding.
+II Related Work
+II-A Radiance Fields and 3D Gaussian Splatting
+Novel view synthesis (NVS) involves rendering unseen viewpoints of a scene from a given set of images.
+One popular approach is Neural Radiance Fields (NeRF), which uses a Multilayer Perceptron (MLP) to represent 3D scenes and map from a 3D coordinate to properties of the scene at the corresponding location.
+It leverages the differentiable volume rendering technique to translate a 3D scene’s continuous representation into 2D images.
+Several works have been proposed to enhance NeRF’s performance by addressing aspects such as speed [78, 79, 10, 59, 24, 25, 72, 94, 60] and adapting it to other tasks [77, 74, 85, 105].
+Though some breakthroughs have been achieved, the reliance on low-efficient volume rendering still hinders real-time rendering.
+Recently, 3D Gaussian Splatting [35] has been proposed as a new technique for novel view synthesis.
+It leverages an ensemble of anisotropic 3D Gaussian splats to represent the scene and employs differentiable splatting for rendering.
+3D Gaussian Splatting has been shown to be an alternative 3D representation of NeRF, benefiting from both its high-quality rendering results and real-time rendering speed.
+Recent research on this technique involves applying 3D Gaussian Splatting to large-scale scene reconstruction [49, 50], the dynamic scenes [52, 87, 46, 112, 98, 1, 51, 75, 31, 45, 95, 41, 80], 3D object and scene generation [81, 93, 16, 89, 90, 111, 43, 103, 109], surface reconstruction [30, 97, 20, 27] and other tasks [9, 86, 44, 75].
+II-B 2D and 3D Perception
+Detection and segmentation are fundamental computer vision tasks.
+Numerous studies [68, 8, 5, 66, 67, 106, 64, 101, 65, 102] have deeply explored various sub-fields.
+Especially, Grounding-DINO [47] enhances the grounding results by introducing the language instructions.
+Traditionally, the segmentation includes three major tasks: semantic [12, 14, 13, 15, 84], instance [28, 48, 4], and panoptic [37] segmentation.
+Due to the operation consistency of the above tasks at the pixel level, many studies have tried to use a unified framework, such as K-net [104], MaskFormer [18], and Mask2Former [17].
+A significant breakthrough in 2D segmentation is the Segment Anything Model (SAM) [38].
+SAM seeks to unify the 2D segmentation task using a prompt-based segmentation approach.
+Its introduction has sparked a new wave of research, with many studies focused on enhancing its functionality.
+These improvements include efficient fine-tuning techniques [70, 71, 34] and distillation-based acceleration methods [107, 99].
+Additionally, SAM has been adapted for various fields, such as medical image analysis [53, 54, 88, 21], concealed object detection [32, 82], image editing [96], remote sensing [11], and 3D bird’s eye view (BEV) sensing [100].
+II-C Segmentation in Radiance Fields
+Neural Radiance Fields (NeRFs) are a popular way of implicitly representing 3D scenes with neural networks.
+Many researchers have explored how to segment objects in 3D using NeRFs for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding.
 Some methods, such as Semantic-NeRF [108], NVOS [73], and SA3D [7], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks.
@@
 Figure 2: Pipeline of our proposed method.
-(a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
-3 Method
-In this section, we first present the preliminary in 3D Gaussian Splatting in Sec.
-3.1 for clear understanding and then define the problem and task in Sec.
-3.2.
+(a) Given a set of clicked points on the $1^{st}$ reference view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
+III Method
+In this section, we first present the preliminary in 3D Gaussian Splatting and Segment Anything Model (SAM) in Sec.
+III-A for clear understanding and then define the problem and task in Sec.
+III-B.
 Our basic training-free pipeline consists of Sec.
-3.3 and Sec.
-3.5.
+III-C and Sec.
+III-E.
 The details of Gaussian Decomposition are described in Sec.
-3.4.
+III-D.
 The pipeline of our method can be seen in Fig.
 2.
-3.1 Preliminary: 3D Gaussian Splatting
+III-A Preliminary
+3D Gaussian Splatting
 3D Gaussian Splatting (3D GS) [35] is an emerging method for real-time radiance field rendering.
@@
 3D GS represents scenes with a set of 3D Gaussians.
-Specifically, each 3D Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma$, an opacity value $\alpha$, and spherical harmonics (SH).
+Specifically, each 3D Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma$ consisting of a scaling factor $s\in\mathbb{R}^{3}$, and a rotation quaternion $q\in\mathbb{R}^{4}$, an opacity value $\alpha$, and spherical harmonics (SH).
+Each 3D Gaussian is characterized by:
+$G(x)=\frac{1}{\left(2\pi\right)^{3/2}\left|\Sigma\right|^{1/2}}e^{-\frac{1}{2}{\left(x-\mu\right)}^{T}\Sigma^{-1}\left(x-\mu\right)}$ (1)
 To render an image, it uses the splatting rendering pipeline, where 3D Gaussians are projected onto the 2D image plane.
@@
 Empirically, for an axis of the ellipse, its length is set to $3\sigma$, where $\sigma$ is the square root the of variance in the axis.
-3.2 Problem Definition
+Segment Anything Model (SAM)
+SAM[38] takes an image $I$ and a set of prompts $P$ as input to output the corresponding segmentation mask $M_{SAM}$:
+$M_{SAM}=SAM(I,P)$ (3)
+where the prompts $P$ can be points, bounding boxes, or texts.
+III-B Problem Definition
 Given a set of trained 3D gaussians $\mathbb{G}=\{\mathbf{g}_{0},\mathbf{g}_{1},\cdots,\mathbf{g}_{n}\}$ and a random initial view $\mathbf{v}_{0}$, users could offer 2D point prompt set $\mathbb{P}_{2D}=\{\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{m}\}$ to specify a 2D object in view $\mathbf{v}_{0}$.
@@
 Thus, our algorithm is designed to minimize the difference between $\mathbf{m}_{i}$ and $\mathbf{m}_{i}^{*}$.
-3.3 Segment 3D Gaussians with 2D Mask
+III-C Segment 3D Gaussians with 2D Mask
 3D Prompts for Multiview Masks Generation.
 By the definition in Sec.
-3.2, users are given the first rendered view to specify the target object.
+III-B, users are given the first rendered view to specify the target object.
 However, only the first view is far from sufficient to segment the target object in 3D space.
@@
 Denoting the $i$-th 2D prompt point in the first given view $\mathbf{v}_{0}$ as $\mathbf{p}_{i}^{0}$, we define the corresponding 3D prompt $\mathbf{p}_{i}^{3D}$ as:
-$\operatorname*{arg\,min}_{\mu}\{d(\mu),d(\mu)>0\mid\mu\in\mathbb{G},\|\mathbf{P}_{0}\mu-\mathbf{p}_{i}^{0}\|_{1}<\epsilon\},$ (2)
+$\arg\min_{\mu}\{d(\mu),d(\mu)>0\mid\mu\in\mathbb{G},\|\mathbf{P}_{0}\mu-\mathbf{p}_{i}^{0}\|_{1}<\epsilon\},$ (4)
 where $d(\mu)$ is the depth of Gaussian center $\mu$ and $\mathbf{P}_{0}$ is the projection for initial view $\mathbf{v}_{0}$.
@@
 This center meets two requirements: (1) it has a similar projected position with $\mathbf{p}_{i}$ with a Manhattan distance less than $\epsilon$, and (2) if there are multiple 3D Gaussian centers satisfying the first requirement, the one with the smallest positive depth is selected as the 3D prompt.
-For all the 2D prompt points in the first view, we could get a set of 3D prompt by Eq.
-2.
+For all the 2D prompt points in the first view, we could get a set of 3D prompts by Eq.
+4.
 Then for another view $\mathbf{v}_{i}$, we project these 3D prompts into the 2D plane, resulting in 2D prompts in the view $\mathbf{v}_{i}$.
@@
 where $\mu_{i}$ is the $i$-th Gausssian center of the scene and $\mathbf{m}_{j}$ is the foreground mask in $j$-th view. $\mathbf{P}_{j}$ stands for the projection matrix of $j$-th view.
-Figure 3: Illustration of the Gaussian Decomposition process.
-3.4 Gaussian Decomposition
+Figure 3: Illustration of the Gaussian Decomposition process. It involves two basic steps: first, to find out the boundary Gaussians and then decompose these Gaussians.
+III-D Gaussian Decomposition
 After obtaining the prompt points of each view, we could assign labels to 3D Gaussians by the projected position of its center.
@@
 However, such a solution greatly damages the 3D structures of the object.
-[…]
+To address this issue, we propose the Gaussian Decomposition to mitigate the boundary roughness while maintaining the 3D structures as complete as possible.
+Fig.
+3 illustrates the basic idea of Gaussian decomposition.
+It has two basic steps to achieve the Gaussian decomposition:
+(1) For each 3D Gaussian, we obtain the corresponding 2D Gaussian by projection and mark it as a boundary Gaussian if one of its long-axis endpoints is outside of the 2D mask.
+Then, as shown in Fig.
+3, we denote the two ends of the long axis of a 2D Gaussian as $A^{*}$ and $B^{*}$, and the intersection of $A^{*}B^{*}$ and mask boundary is $O^{*}$.
+$A$, $B$, and $O$ are the corresponding 3D points in the long axis of the 3D Gaussian.
+Assuming $A^{*}$ is in the mask while $B^{*}$ is out of the mask, we define
 $\lambda_{3D}=\frac{OA}{AB},$ (6)
@@
 8 indicates that Gaussian Splatting simplifies the perspective projection to an affine projection, thus decomposition scaling ratio $\lambda_{3D}$ in 3D space is equivalent to the ratio $\lambda_{2D}$ in the 2D plane.
-[…]
+Let $\mathbf{g}$ denote a 3D Gaussian across the boundary. Its scale in the long axis and the 3D center are defined as $s$ and $\mathbf{\mu}$, respectively. We have
 $s^{\prime}=\lambda_{2D}s,$ (9)
@@
 Another decomposed Gaussian outside of the mask is removed.
-3.5 Multiview Label Voting
+III-E Multiview Label Voting
 So far, every 3D Gaussian including the decomposed one has a list of binary labels $\mathbf{L}_{i}$, using the label assignment in Sec.
-3.3.
+III-C.
 Leveraging the assigned labels, here we adopt a simple yet effective heuristic rule to determine if a 3D Gaussian $\mathbf{g}_{i}$ belongs to the target 3D object.
@@
 $s_{i}=\frac{1}{N}\sum_{j=0}^{N-1}L_{ij},$ (11)
-where $N$ is the number of views. Then we adopt a threshold $\tau$ and regard a Gaussian with a score higher than $\tau$ as positive.
-4 Experimets
-4.1 Datasets
+where $N$ is the number of views.
+First, $s_{i}>0.5$ is performed to vote out the object 3D Gaussian.
+Then, to reduce the background bias due to occlusion, we adopt a threshold $\tau$ and an Object-ID $\mathbf{O}_{i}$ for each 3D Gaussian $\mathbf{g}_{i}$ is determined by:
+$\mathbf{O}_{i}=\begin{cases}1&\text{if }s_{i}>\tau,\\ 0&\text{if }s_{i}<\tau,\end{cases}$ (12)
+III-F Multi-object Segmentation
+There are numerous objects in the 3D scenes.
+Our method can also support segmenting all objects simultaneously by merely extending the binary masks to multi-label masks.
+Specifically, the element $\mathbf{L}_{ij}$ in the matrix $\mathbf{L}$ is extended from binary values to multiple values, where $\mathbf{L}_{ij}\in\{0,1,...,C-1\}$ and $C$ is the number of objects.
+The similar multi-view label voting strategy will determine an Object-ID $\mathbf{O}_{i}$ for each 3D Gaussian $\mathbf{g}_{i}$:
+$\mathbf{O}_{i}=Mode(\mathbf{L}_{ij}),j\in\{0,1,...,N-1\},$ (13)
+where $Mode(.)$ is the function of finding the element that appears the most, and $N$ is the number of views.
+IV Experimets
+IV-A Datasets
 We choose different datasets to testify our method, including LLFF [57], Mip-NeRF 360 [3], LERF [36], and some test scenes from the 3D Gaussian Splatting [35].
 These datasets contain both small indoor objects and large outdoor scenes, which are very complex and challenging.
-For quantitative experiments, because there is no existing benchmark that can be used in 3D Gaussian space, we use the SPIn-NeRF [23] dataset with 2D ground-truth for evaluation.
-4.2 Quantitative results
-Table 1: Quantitative results on SPIn-NeRF dataset.
-“Singe view” denotes projecting the 2D segmentation result to 3D simply, thus we don’t consider its time cost.
-There is extra time cost (about 10mins) for SAGA to distill 2D knowledge to 3D-GS, while other methods don’t need this process.
+For quantitative experiments, because there is no existing benchmark that can be used in 3D Gaussian space, we use the SPIn-NeRF [58] dataset with 2D ground truth for evaluation.
+IV-B Implementation Details
+As the implementation of 3D Gaussian Splatting [35] for each scene, we follow the official code with default parameters, and each scene is trained with 30000 iterations.
+For generating single object segmentation masks, we don’t limit the number of clicked points on a single reference view for SAM.
+Binary masks of other views can be automatically generated via the proposed method.
+This is reasonable since users can refine their input prompts to help SAM generate a 2D mask as accurately as possible from the reference view.
+For scene segmentation to generate multi-label masks, we first employ SAM to produce instance masks for individual views under automatic segmentation mode.
+To ensure 2D mask consistency across views, we use a pre-trained zero-shot tracker [19, 92] to propagate and associate masks.
+For the selection of hyper-parameters, we only have one to control, which is the confidence score threshold $\tau$ in the Multiview Label Voting method III-E.
+By default, we set the value to 0.7 in all experiments.
+Also, in practical use, users can manually set this value according to the complexity of different scenes.
+As for the number of views used in experiments, we follow SA3D [7] to select all view images of each scene to finish our segmentation process.
+IV-C Quantitative results
+TABLE I: Quantitative results on SPIn-NeRF dataset. Singe view denotes projecting the 2D segmentation result to 3D simply.
 Scenes Single View [7] MVSeg [58] SA3D [7] SAGA [6] Ours
@@
 Horns 85.3 97.1 92.8 98.7 94.5 99.0 - - 91.1 98.4
-Fortress 94.1 99.1 97.7 99.7 98.3 99.8 - - 96.6 99.5
+Fortress 94.1 99.1 97.7 99.7 98.3 99.8 - - 95.6 99.5
 Fork 69.4 98.5 87.9 99.5 89.4 99.6 - - 83.4 99.3
-Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 - - 92.0 99.0
+Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 - - 92.6 99.0
 Truck 37.9 77.9 85.2 95.1 90.8 96.7 - - 93.0 97.9
-Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 88.4 99.7
-Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 89.9 98.7
-Time cost - 3-6mins 2-10mins 0.08s-0.9s within 15s
-(/object) (+10mins)
-We first conduct experiments on the SPIn-NeRF [23] dataset for quantitative analysis.
-Given a set of images of the scene, we follow the process described in Section 3 to obtain the segmentation of the target object in 3D Gaussian Space.
+Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 90.2 99.7
+Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 90.0 98.7
+Point-Guided Segmentation: We first conduct experiments on the SPIn-NeRF [58] dataset for quantitative analysis.
+Given a set of images of the scene, we follow the process described in Section III to obtain the segmentation of the target object in 3D Gaussian Space.
 The segmented 3D Gaussians are used to render 2D masks in other views.
 Finally, we calculate the IoU and Accuracy between these rendered and the ground-truth masks.
-Results can be seen in Table 1.
+Results can be seen in Table I.
 In the comparison, it is noteworthy that both the MVSeg [58] and the SA3D [7] require additional parameters and a computation-costly training process.
@@
 In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg.
-When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.8% IoU and +3.5% Acc.
+When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.9% IoU and +3.5% Acc.
 These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
-Table 2: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
+TABLE II: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
 Scene room truck fortress pinecone
@@
 Acc 97.9 97.8 98.8 97.0
+Text-Guided Segmentation: We replace the point prompts with text prompts corresponding to the objects.
+Given a text input, we prompt the Grounding DINO [47] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks.
+Then, following the segmentation process in Section III, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input.
+Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks.
+Quantitative results are shown in Table II.
+For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts.
+These results demonstrate our method can combine multi-modal prompts as input.
+IV-D Boundary Segmentation Analysis
+The main contribution of our work is to address the boundary issues of segmentation in 3DGS, as mentioned in Sec.
+I and Fig.
+1.
+To validate the effectiveness of our work, we evaluate metrics, including IoU, AP, and F1-score, on region nearby boundaries.
+Specifically, we extract parts of 3-pixel width alongside boundaries of ground-truth masks and compare results with previous work SAGA [6].
+Experiments are conducted on Scene pinecone and fortress from the LLFF [57] dataset.
+Better results of ours are indicated in Tab.
+III.
+Benefiting from our proposed Gaussian Decomposition for ambiguous geometry around boundaries, we can achieve more complete and accurate boundary segmentation results.
 Figure 4: Qualitative results compared with SA3D [7] and SAGA [6] in different scenes (LERF-figurines [36], SPIn-NeRF-Orchids [58], LERF-dozer-nerfgun-waldo [36]).
 We enlarge the boxed area on the right for a better visualization.
-Next, we replace the above point prompts with text prompts corresponding to the objects.
-Given a text input, we prompt the Grounding DINO [18] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks.
-Then, following the segmentation process in Section 3, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input.
-Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks.
-Quantitative results are shown in Table 2.
-For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts.
-These results demonstrate our method can combine multi-modal prompts as input.
-4.3 Qualitative results
+TABLE III: Evaluation on boundary masks. We extract parts of 3-pixel width alongside boundaries of ground truth and segmentation masks and evaluate on IoU, AP, and F1-score metrics.
+Scene IoU AP F1-score
+SAGA [6] Ours SAGA Ours SAGA Ours
+pinecone 13.8 21.8 15.7 23.5 24.0 35.6
+fortress 16.7 23.3 31.8 60.2 28.5 37.2
+IV-E Qualitative results
 We conduct four kinds of tasks to demonstrate the potential application and qualitative performance of our approach, including point-guided segmentation, text-guided segmentation, scene editing, and collision detection.
-Figure 5: 3D segmentation results with text prompts in Mip-NeRF360-garden [2].
+Figure 5: 3D segmentation with text prompts in Mip-NeRF360-garden [3].
 Point-Guided Segmentation: We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image.
@@
 Text-Guided Segmentation: We further conduct 3D object segmentation experiments guided by text prompts.
-We follow the same process as in Table 2 experiments to conduct visualization experiments.
+We follow the same process as in Table II experiments to conduct visualization experiments.
 In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [3], using three text prompts including “The table”, “The vase” and “The bonsai”.
@@
 Comparison between 2D and 3D Segmentation: Our method improves the segmentation quality of SAM.
-Due to the sensitivity of SAM to the scene viewpoint, the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig.
-6).
-Table 3: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset.
+Due to the sensitivity of SAM to the scene viewpoint (e.g.
+illumination), the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig.
+7, $1^{st}$ Row).
+Furthermore, we can obtain detailed 3D segmentation results under noisy predictions from SAM.
+As shown in Fig.
+7, $2^{nd}$ Row, the multi-view voting ensures that object Gaussians are retained correctly and background Gaussians obtained from the noisy boundaries of SAM predictions are removed.
+Scene Editing: Scene editing is a basic application for 3D reconstructed scenes.
+However, it’s quite difficult to do this without being able to locate the specific objects.
+This task demonstrates the ability of our approach to help edit the 3D scenes.
+Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them.
+Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Fig.
+6.
+The instances in 1st column with red bounding boxes are objects to be segmented.
+It can be seen that with objects segmented in 3D Gaussians, they can be translated and rotated in any direction in the scene.
+After the removal of segmented objects, original scenes can still keep intact.
+Collision Detection: Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators.
+In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space.
+We choose two scenes from SPIn-NeRF dataset.
+Following the process of our segmentation method, we can obtain the corresponding segmented 3D Gaussian points in the scene.
+To this end, we use the Quickhull [2] algorithm to build the collision mesh upon our segmented objects.
+The results in the last column of Fig.
+6 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
+Figure 6: Visualization examples of scene editing after the object segmentation.
+We offer three scene editing examples: removal, translation, and rotation.
+We further provide collision mesh computed after segmentation, as shown in the last column.
+IV-F Time Cost Analysis
+We conduct the time cost analysis compared with previous 3DGS segmentation methods, namely SAGA [6] and Gaussian Grouping [92].
+We select the figurines scene from the LERF dataset [36], which includes 20+ objects, using a single NVIDIA A100 GPU for fair evaluation.
+All methods utilize SAM to extract image features and then perform mask generation.
+Thus, they share a similar time cost, roughly 2 minutes to process 300 images (SAM Extraction in Tab.
+IV).
+This procedure can be regarded as a pre-processing and can be accelerated by image-batching or deploying Efficient-SAM.
+For extra processing time, both baseline methods require gradient descent optimization over 30,000 iterations to distill SAM features [6] or ID embedding [92] into 3D space linked with each 3D Gaussian, resulting in considerable additional training time for re-optimizing 3D scenes.
+On the contrary, our method only needs to perform multi-view and multi-label assignments, the time cost of which is merely 12 seconds.
+As for segmenting a single object, our method merely requires an efficient voting process, taking just 0.4 milliseconds, which is much faster than other methods as they both demand network forwarding.
+Figure 7: Comparison between 2D and 3D Segmentation. Noisy predictions exist in SAM’s segmentation, such as incomplete objects and background regions. These issues are solved in our 3D segmentation.
+TABLE IV: Time cost comparisions on figurines scene with 20+ objects to be segmented simultaneously.
+Time SAM Extra Optimization Forward
+Extraction Processing Steps Segmentation
+SAGA [6] 2 min 10 min 30000 0.5s
+Ga-Grouping [92] 9 min 30000 0.3s
+Ours 12s 1 0.4ms
+TABLE V: Ablation on Gaussian decomposition (GD) on SPIn-NeRF dataset.
 The 2nd and 3rd columns compare the results before and after using the Gaussian decomposition (GD).
@@
 Fork 83.4 99.3 81.8 99.2 79.9 99.1
-Pinecone 92.1 98.9 91.6 98.9 82.9 97.5
-Truck 94.0 97.9 93.4 97.8 91.4 96.8
-Lego 88.4 99.7 88.4 99.6 82.9 99.4
-mean 89.9 98.7 86.6 98.1 84.2 97.7
-Scene Editing: Scene editing is a basic application for 3D reconstructed scenes.
-However, it’s quite difficult to do this without being able to locate the specific objects.
-This task demonstrates the ability of our approach to help edit the 3D scenes.
-Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them.
-Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Fig.
-7.
-The instances in 1st column with red bounding boxes are objects to be segmented.
-It can be seen that with objects segmented in 3D Gaussians, they can be translated and rotated in any direction in the scene.
-After the removal of segmented objects, original scenes can still keep intact.
-Figure 6: Comparison between 2D and 3D Segmentation.
-Collision Detection: Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators.
-In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space.
-We choose two scenes from SPIn-NeRF dataset.
-Following the process of our segmentation method, we can obtain the corresponding segmented 3D Gaussian points in the scene.
-To this end, we use the Quickhull [1] algorithm to build the collision mesh upon our segmented objects.
-The results in the last column of Fig.
-7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
-Figure 7: Visualization examples of scene editing after the object segmentation.
-We offer three scene editing examples: removal, translation, and rotation.
-We further provide collision mesh computed after segmentation, as shown in the last column.
-4.4 Ablation Study
-Gaussian Decomposition:
-Gaussian Decomposition is proposed to address the issue of roughness boundaries of 3D segmented objects, which results from the non-negligible spatial sizes of 3D Gaussian located at the boundary.
+Pinecone 92.6 98.9 91.6 98.9 82.9 97.5
+Truck 93.0 97.9 93.4 97.8 91.4 96.8
+Lego 90.2 99.7 88.4 99.6 82.9 99.4
+mean 90.0 98.7 86.6 98.1 84.2 97.7
+IV-G Ablation Study
+Gaussian Decomposition: Gaussian Decomposition is proposed to address the issue of roughness boundaries of 3D segmented objects, which results from the non-negligible spatial sizes of 3D Gaussian located at the boundary.
 We conduct both quantitative and qualitative experiments to verify the effectiveness of our approach.
-For quantitative ablation experiments, Table 3 shows the results.
+For quantitative ablation experiments, Table V shows the results.
 We compare our proposed Gaussian Decomposition scheme with two other processing methods, one for segmentation without any special handling (the 3rd column), and the other for directly removing such Gaussians (the last column).
-The experiments are conducted on SPIn-NeRF dataset, following the same evaluation process as Table 1.
-Compared with the others, our proposed approach outperforms them on all scenes with +3.3% and +5.7% IoU, respectively, demonstrating the effectiveness of Gaussian Decomposition.
+The experiments are conducted on SPIn-NeRF dataset, following the same evaluation process as Table I.
+Compared with the others, our proposed approach outperforms them on all scenes with +3.4% and +5.8% IoU, respectively, demonstrating the effectiveness of Gaussian Decomposition.
 Besides, it can seen that directly removing these gaussians cannot decrease the roughness of mask boundaries, leading to even 2.4% lower IoU values.
@@
 9 shows the visualization results for comparison.
-We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column).
+We select two representative scenes in which large-scale Gaussian points exist at the junction/contact (can be seen in the 2nd column).
 This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy.
 Though this idea is simple, the improvement is obvious.
-Table 4: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
+TABLE VI: Ablation on different number of views for 3D segmentation. Numbers in parentheses represent the used view percentage of the total training view.
 Number of views 5(10%) 10(20%) 21(50%) 42(100%)
-IoU on Fortress 91.16 92.11 93.82 96.55
+IoU on Fortress 91.16 92.11 93.82 95.6
 Number of views 25(10%) 50(20%) 125(50%) 251(100%)
-IoU on truck 90.27 90.97 92.11 93.48
-Table 5: Ablation on confidence score threshold $\tau$ on the “truck” scene.
-$\tau$ 0.3 0.5 0.6 0.7 0.9
-IoU 61.7 82.7 93.4 90.6 53.9
-Selection of Views:
-In this section, we study the influence of selected 2D views on 3D Gaussian segmentation.
+IoU on truck 90.27 90.97 92.11 93.0
+TABLE VII: Ablation on confidence score threshold $\tau$ on the truck scene.
+Scene / $\tau$ 0.6 0.65 0.7 0.8
+pinecone 90.8 92.0 92.6 91.4
+fortress 94.0 95.4 95.6 96.2
+lego 89.7 89.8 90.2 90.2
+Figure 8: Ablation on view selection, including view directions and the number of views.
+Selection of Views: In this section, we study the influence of selected 2D views on 3D Gaussian segmentation.
 Fig.
@@
 Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views.
-Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase.
+Table VI also draws a similar conclusion: as the number of views increases, the IoU values will also increase.
 It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach.
@@
 This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
-Hyper-Parameters:
-In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec.
-3.5.
-As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects.
-For scenes with some views that may cover only part of the objects to be segmented (i.e.
-half of the truck body in the “truck” scene), a small $\tau$ value is preferred.
-In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
-5 Discussions and Limitations
+Hyper-Parameters: In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec.
+III-E.
+To evaluate the robustness of the selection of the hyper-parameter under different scenes, we set different values to conduct experiments on Scene pinecone, fortress, and lego from LLFF [57] dataset.
+As indicated in Tab.
+VII, our method is not sensitive to the threshold selection.
+Robust results are observed in multiple scenes when setting the threshold between 0.6 and 0.8.
+Values within this range can obtain fine results.
+In all of our experiments, the value of the hyper-parameter $\tau$ is set to 0.7 by default.
+V Discussions and Limitations
 Through experiments, we found that our method does not perform well in objects where 3D Gaussians are very sparse, such as the LLFF-room [57] scene.
@@
 We believe this limitation can be alleviated by future research in a more structured 3D-GS representation, yielding more accurate results.
-Figure 9: Ablation on view selection.
-6 Conclusion
+VI Conclusion
 We address the issue that rough/incomplete boundaries in segmenting 3D-GS and propose a novel Boundary-enhanced segmentation pipeline.